ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: Dashkova, O. Yu. , On groups with restrictions on non-normal subgroups Locally nilpotent groups in which the finiteness condition of special rank is imposed on the system of all non-normal subgroups of group are studied. Tte structure and properties of such groups are described.